# GroundCap: A Visually Grounded Image Captioning Dataset

### v2

@@ Abstract
Current image captioning systems lack the ability to link descriptive text to specific visual elements, making their outputs difficult to verify. While recent approaches offer some grounding capabilities, they cannot track object identities across multiple references or ground both actions and objects simultaneously. We propose a novel ID-based grounding system that enables consistent object reference tracking and action-object linking, and present GroundCap, a dataset containing 52,016 images from 77 movies, with 344 human-annotated and 52,016 automatically generated captions. Each caption is grounded on detected objects (132 classes) and actions (51 classes) using a tag system that maintains object identity while linking actions to the corresponding objects. Our approach features persistent object IDs for reference tracking, explicit action-object linking, and segmentation of background elements through K-means clustering. We propose gMETEOR, a metric combining caption quality with grounding accuracy, and establish baseline performance by fine-tuning Pixtral-12B. Human evaluation demonstrates our approach’s effectiveness in producing verifiable descriptions with coherent object references.
@@ 1 Introduction
One of the primary goals combining computer vision and natural language processing is to enable machines to understand and communicate about visual scenes. This objective encompasses numerous tasks, including recognizing objects, describing their attributes and relationships, and providing contextually relevant descriptions of scenes [1]. While significant progress has been made in image classification, object detection, and image captioning, a critical aspect of human visual communication remains under-explored: the ability to ground language to specific elements within an image.
@@
We introduce GroundCap, a novel dataset and task formulation that addresses the challenge of visually grounded image captioning. GroundCap is designed to bridge this gap by providing a large-scale dataset of images with captions that are explicitly grounded to visual elements, mimicking the human ability to reference objects when describing a scene.
GroundCap comprises 52,016 images extracted from 77 movies from the MovieNet dataset [2]. We consider two caption types: 52,016 automatically generated captions (42,016 for training and 10,000 for evaluation) and 344 human-annotated captions (274 for training and 70 for evaluation). Each image is further annotated with object positions and classes from a set of 132 object classes. The captions are grounded, using a novel tag format that links textual descriptions to multiple visual elements. These captions not only reference objects but also describe actions, and the actions themselves are grounded on one or more objects, providing for rich contextual descriptions.
In contrast with works that focus solely on object grounding [3, 4], GroundCap introduces several key innovations. Our framework assigns unique identifiers to each object and unifies the grounding of objects, actions, and locations while maintaining consistent object identity throughout descriptions. This approach enables rich contextual descriptions that can reference multiple objects from different classes within a single text span and resolve complex references (e.g., linking pronouns like “They” to specific objects like “person-1” and “person-2”). Our system can ground spans of text to specific portions of background elements, moving beyond treating them as single entities.
Our main contributions can be summarized as follows: (1) a novel dataset of 52,016 images with grounded captions that maintain object identity across references; (2) a unified grounding framework that links objects, actions, and locations while handling background elements; (3) gMETEOR, a new metric that combines caption quality with grounding accuracy; (4) a strong baseline model based on fine-tuned Pixtral-12B; (5) comprehensive evaluation through automated metrics, human assessment, and ChatGPT-4o evaluation; The dataset and fine-tuned model are available on Hugging Face.
The paper is organized as follows: Section 2 reviews related work on image captioning, object detection, and grounded captioning. Section 3 details the creation of GroundCap. Section 4 introduces gMETEOR, a novel metric designed to assess the quality of grounded captions. Section 5 presents a baseline model for GroundCap. Section 6 outlines the results of human evaluation of GroundCap, baseline model and human-refined captions. Section 7 outlines current limitations. Section 8 presents the conclusions and future work.
@@ 2.2 Image Captioning
Image captioning datasets combine visual and textual information, requiring models to generate descriptive sentences for images. Flickr30k [12] and MS COCO Captions [1] are widely used benchmarks in this domain. These datasets, however, typically provide only general descriptions without explicit grounding to specific image regions or objects. Visual Genome [13] takes a step towards grounded captioning by providing several region-specific captions for each image, where each caption describes a particular area defined by a single bounding box. While this approach offers more localized descriptions, it differs from our task in GroundCap. Recent advances in Multimodal Large Language Models, such as Flamingo [14], LLaVA [15] and Pixtral-12B [16] have significantly improved image captioning capabilities. While some of these models, like Qwen-VL, provide visual grounding capabilities, they lack the ability to maintain consistent object identities across multiple references within a caption.
2.3 Action Recognition
Datasets for action recognition in images and videos, such as HICO-DET [17] and AVA [18], are relevant as they annotate human actions and interactions with objects. HICO-DET provides annotations for 600 human-object interaction categories in static images, while AVA offers dense annotations of 80 atomic visual actions in movie clips. IXMAS [19] uses multiple camera views to record actions, enabling viewpoint-independent action recognition. HMDB [20] provides a large collection of 51 human action categories with over 6,766 video clips extracted from various sources. The UCF dataset [21] specifically addresses the challenge of recognizing actions “in the wild”, featuring videos that capture real-world scenarios.
2.4 Grounded Captioning
Several works have explored aspects of grounded captioning, focusing on linking textual descriptions to visual elements.
Visual Genome [13], provides multiple region descriptions per image, each associated with a single bounding box. This approach limits the ability to create dense descriptions where words can be simultaneously grounded to multiple objects and actions across regions if the frame. The region-specific approach also constrains the ability to capture scene descriptions that integrate both static elements and dynamic interactions within a caption.
@@ 2.4 Grounded Captioning
Recent work has explored different approaches to visual grounding using Large Language Models (LLMs). GROUNDHOG [4] uses pixel-level segmentation with unified <GRD> tags, Groma [22] employs region tokens for arbitrary region grounding, and KOSMOS-2 [3] represents coordinates as discrete location tokens in a Markdown-like format. Florence-2 [23] performs grounding through a sequence-to-sequence architecture that quantizes coordinates into 1,000 bins and represents regions in various formats (box, quad box, or polygon) depending on the task requirements. Panoptic Narrative Grounding (PNG) [24] grounds descriptions to panoptic segmentation regions using mouse trace annotations and WordNet-based [25] semantic matching. [26] proposes a one-stage weakly-supervised approach that processes raw RGB images and computes visual language attention maps.
Despite their promising performance, these models lack the ability to track object identities across multiple references or ground both actions and objects simultaneously.
@@ 3 The GroundCap Dataset
This section details the creation of GroundCap [^1] The GroundCap dataset is publicly available at https://huggingface.co/datasets/daniel3303/GroundCap.. We first describe our frame selection process from MovieNet, followed by the object detection approach that handles both distinct objects and background elements. We then explain the caption generation pipeline that combines automated generation with human refinement.
@@ 3.2.2 Bounding Box Generation for “Stuff” Classes
Our approach to generating bounding boxes for “stuff” classes aims to provide a detailed representation of background elements while balancing with the lowest number of bounding boxes possible. The process begins by extracting pixel coordinates of the segmented area. K-means clustering is then applied to these coordinates, with K ranging from 1 to 6. For each K, we perform 10 initializations to mitigate the impact of the sensitivity of K-means to initial conditions. This strategy increases the likelihood of finding a globally optimal solution.
@@
For each clustering result, we use a systematic method to generate initial bounding boxes. We aim to capture the core of each cluster while being resilient to outliers. Specifically, we calculate the 5th and 95th percentiles of the $x$ and $y$ coordinates within each cluster. These percentiles form the basis of our initial bounding boxes, effectively excluding the 5% most extreme points on each end. This approach strikes a balance between coverage and compactness, ensuring that the initial boxes are representative of the cluster’s main body without being unduly influenced by scattered outlier points.
@@
After generating these initial boxes, we perform an adjustment step to eliminate overlaps between boxes of the same class while striving to maintain the integrity of each box’s representation. This adjustment is necessary because the independent generation of boxes for each cluster can lead to overlaps, especially in areas where clusters are closely spaced or intermingled. Our adjustment algorithm iteratively identifies overlapping pairs of boxes and modifies their boundaries. When adjusting for overlaps, we prioritize minimal distortion by modifying the overlapping edges proportionally to the size of each box. We prefer adjusting the lowest overlapping axis with respect to the length of the boxes in the corresponding axis.
@@
For instance, we compare the horizontal overlap relative to the total width of the overlapping boxes against the vertical overlap relative to their total height. If the proportional horizontal overlap is smaller, we adjust the left and right edges; otherwise, we adjust the top and bottom edges, thus preserving the aspect ratio of the boxes as much as possible. The adjustment for each box is calculated based on its size relative to the total size of both overlapping boxes in the chosen dimension (i.e., larger boxes undergo larger adjustments). The process results in a set of non-overlapping boxes that closely adhere to the underlying cluster distributions. It provides a clean and accurate representation of the “stuff” class regions, balancing the need for coverage of the segmented area with the desire for minimal overflow into non-segmented regions and the goal of minimizing the number of bounding boxes. The final set of boxes is evaluated based on two primary criteria.
  • 1.
    Coverage: The proportion of the segmented area that is enclosed by the bounding boxes. We aim for a minimum coverage of 90% to ensure that the majority of the “stuff” region is represented.
  • 2.
@@
    Overflow: The proportion of the area within the bounding boxes that does not belong to the segmented region. We aim for a maximum allowed overflow of 30% into non-segmented areas.
@@
The evaluation function balances these criteria, seeking to maximize coverage while minimizing overflow. This process is repeated for each number of clusters and each initialization until and valid configuration is found. A configuration is considered valid when the coverage exceeds 90% and the overflow is below 30% or when the maximum number of clusters is reached. If multiple valid configurations are found for the same number of clusters, the best one is selected based on the configuration score calculated as the difference between coverage and overflow. This allows us to represent large or disconnected background elements with multiple, focused bounding boxes.
@@ 3.2.3 Post-processing
We assign unique IDs to each detected object, following their spatial sorting order: e.g. person-0 will always appear before person-1 in the image, either in a higher band (closer to the top) or in the same band but further to the left. The consistent ordering ensures that object references in captions are predictable and spatially meaningful – when referring to “the person in the top left corner”, we can be confident it corresponds to person-0.
@@ 3.3 Grounded Caption Generation
The process begins by generating a general caption for the main image using Pixtral [16]. This initial caption provides a high-level description of the scene without explicit grounding information. The model receives the image along with a system prompt explaining the captioning task and a set of few-shot examples demonstrating proper caption format and style, following the few-shot learning paradigm established in recent work [15]. We then generate focused captions for each detected object. These object-specific captions are created by providing the model with cropped regions corresponding to the bounding box of each object. In the final stage, the LLM synthesizes a grounded caption by combining the main caption, object-specific captions, and detection information, such as object metadata, including class, ID, and coordinates of the bounding box, and a list with 51 allowed actions taken from HMDB [20]. The model generates a grounded caption incorporating our grounding tag system, using three distinct tags: <gdo> for objects, <gda> for actions, and <gdl> for locations.
Quality control is implemented through an evaluation system based on precision ($P$), recall ($R$), and F1 score metrics. These values follow standard information retrieval definitions: $P=\text{TP}/(\text{TP}+\text{FP})$; $R=\text{TP}/(\text{TP}+\text{FN})$; and F1 score as the harmonic mean of precision and recall. TP (true positives) represents correctly grounded objects, FP (false positives) are objects mentioned in the caption but not present in detections, and FN (false negatives) are detected objects omitted from the caption.
@@ 4 Grounded METEOR (gMETEOR)
5 Baseline Model
To establish a baseline for grounded captioning on our dataset, we fine-tuned Pixtral-12B [16] on GroundCap  [^2] Fine-tuned model: https://huggingface.co/daniel3303/PixtralGroundCap.. We chose Pixtral-12B for its strong performance on vision-language tasks when compared with models from similar size.
Our baseline processes the full image along with metadata about each detected object (including object ID, class, and normalized coordinates), enabling it to generate captions with appropriate grounding tags (<gdo>, <gda>, and <gdl>) that link text spans to specific objects in the image.

### v3

@@ Abstract
Current image captioning systems lack the ability to link descriptive text to specific visual elements, making their outputs difficult to verify. While recent approaches offer some grounding capabilities, they cannot track object identities across multiple references or ground both actions and objects simultaneously. We propose a novel ID-based grounding system that enables consistent object reference tracking and action-object linking. We present GroundCap, a dataset containing 52,016 images from 77 movies, with 344 human-annotated and 52,016 automatically generated captions. Each caption is grounded on detected objects (132 classes) and actions (51 classes) using a tag system that maintains object identity while linking actions to the corresponding objects. Our approach features persistent object IDs for reference tracking, explicit action-object linking, and the segmentation of background elements through K-means clustering. We propose gMETEOR, a metric combining caption quality with grounding accuracy, and establish baseline performance by fine-tuning Pixtral-12B and Qwen2.5-VL 7B on GroundCap. Human evaluation demonstrates our approach’s effectiveness in producing verifiable descriptions with coherent object references.
@@
{graphicalabstract}
GroundCap: A Visually Grounded Image Captioning Dataset Input[Uncaptioned image] Movie Frames from MovieNet
Selected for rich visual content
Detection[Uncaptioned image] 132 Object Classes 50 Action Classes ID-based Tracking + K-means Clustering Generation[Uncaptioned image] Object (GDO) Action (GDA) Location (GDL) Multi-level Grounding FramesObjects 52,016 77 Captions Movies 344 Human-annotated + 52,016 Auto-generated Evaluation: gMETEOR Metric + Human Assessment Baseline: Fine-tuned Pixtral-12B Open-Source dataset & Open Weight model
{highlights}
Novel dataset GroundCap features 52,016 images with 52,350 grounded captions.
Captions link actions or multiple objects, describing their interactions and roles.
gMETEOR metric combines caption quality and grounding accuracy to evaluate captions.
Fine-tuned Pixtral-12B and Qwen2.5VL-7B on GroundCap set strong captioning baselines.
Human evaluation studies validate the quality of the grounded captions.
@@ 1 Introduction
One of the primary goals of combining computer vision and natural language processing is to enable machines to understand and communicate about visual scenes. This objective encompasses numerous tasks, including recognizing objects, describing their attributes and relationships, and providing contextually relevant descriptions of scenes [1]. While significant progress has been made in image classification, object detection, and image captioning, a critical aspect of human visual communication remains under-explored: the ability to ground language to specific elements within an image.
@@
We introduce GroundCap, a novel dataset and task formulation that addresses the challenge of visually grounded image captioning. GroundCap is designed to bridge this gap by providing a dataset of images with captions that are explicitly grounded to visual elements, mimicking the human ability to reference objects when describing a scene.
GroundCap comprises 52,016 images extracted from 77 movies from the MovieNet dataset [2]. We consider two caption types: 52,016 automatically generated captions (42,016 for training and 10,000 for evaluation) and 344 human-annotated captions (274 for training and 70 for evaluation). Each image is further annotated with object positions and classes from a set of 132 object classes. The captions are grounded using a novel tag format that links textual descriptions to multiple visual elements. These captions not only reference objects but also describe actions, and the actions themselves are grounded on one or more objects, providing for rich contextual descriptions.
In contrast to works that focus solely on object grounding [3, 4], GroundCap introduces several key innovations. Our framework assigns unique identifiers to each object and unifies the grounding of objects, actions, and locations while maintaining consistent object identity throughout descriptions. This approach enables rich contextual descriptions that can reference multiple objects from different classes within a single text span and resolve complex references (e.g., linking pronouns like “They” to specific objects like “person-1” and “person-2”). Our system can ground spans of text to specific portions of background elements, moving beyond treating them as single entities.
Our main contributions can be summarized as follows: (1) a novel dataset of 52,016 images with grounded captions that maintain object identity across references; (2) a unified grounding framework that links objects, actions, and locations while handling background elements; (3) gMETEOR, a new metric that combines caption quality with grounding accuracy; (4) strong baseline models based on fine-tuned Pixtral-12B and Qwen2.5-VL 7B; (5) evaluation through automated metrics, human assessment, and ChatGPT-4o evaluation; The dataset and fine-tuned model are available on Hugging Face.
The paper is organized as follows: Section 2 reviews related work on image captioning, object detection, and grounded captioning. Section 3 details the creation of GroundCap. Section 4 introduces gMETEOR, a novel metric designed to assess the quality of grounded captions. Section 5 presents baseline models for GroundCap. Section 6 outlines the results of human evaluation of GroundCap, baseline model and human-refined captions. Section 7 outlines current limitations. Section 8 presents the conclusions and future work.
@@ 2 Related Work
Image captioning datasets combine visual and textual information, requiring models to generate descriptive sentences for images. Flickr30k [12] and MS COCO Captions [1] are widely used benchmarks in this domain. These datasets, however, typically provide only general descriptions without explicit grounding to specific image regions or objects. Visual Genome [13] takes a step towards grounded captioning by providing several region-specific captions for each image, where each caption describes a particular area defined by a single bounding box. While this approach offers more localized descriptions, it differs from our task in GroundCap. DenseCap [14] introduced dense captioning for localizing and describing salient regions in images. Recent advances in Multimodal Large Language Models, such as Flamingo [15], LLaVA [16], Pixtral-12B [17] and Qwen2.5-VL [18] have significantly improved image captioning capabilities. While some of these models, like Qwen-VL, provide visual grounding capabilities, they lack the ability to maintain consistent object identities across multiple references within a caption.
@@
Several works have explored aspects of grounded captioning, focusing on linking textual descriptions to visual elements. Visual Genome [13], provides multiple region descriptions per image, each associated with a single bounding box. This approach limits the ability to create dense descriptions where words can be simultaneously grounded to multiple objects and actions across regions if the frame. The region-specific approach also constrains the ability to capture scene descriptions that integrate both static elements and dynamic interactions within a caption. Recent work has explored different approaches to visual grounding using Large Language Models (LLMs). GROUNDHOG [4] uses pixel-level segmentation with unified <GRD> tags, Groma [19] employs region tokens for arbitrary region grounding, and KOSMOS-2 [3] represents coordinates as discrete location tokens in a Markdown-like format. Florence-2 [20] performs grounding through a sequence-to-sequence architecture that quantizes coordinates into 1,000 bins and represents regions in various formats (box, quad box, or polygon) depending on the task requirements. Panoptic Narrative Grounding (PNG) [21] grounds descriptions to panoptic segmentation regions using mouse trace annotations and WordNet-based [22] semantic matching. [23] proposes a one-stage weakly-supervised approach that processes raw RGB images and computes visual language attention maps.
Several large-scale grounded datasets have been developed to support these approaches. GLaMM [24] introduced GranD-f, a dataset of 214K image-grounded text pairs. Groma [19] curated Groma Instruct with 30K visually grounded conversations. KOSMOS-2 [3] was trained on GrIT, a dataset of 91M images with grounded image-text pairs, though only a small subset is publicly available. Florence-2 was trained on FLD-5B [20], containing 126M images with 5.4B annotations, but the dataset remains proprietary. GROUNDHOG utilizes the M3G2 dataset with 2.5M images, which is also not publicly available. The availability of training data remains a significant factor in the reproducibility and advancement of grounded captioning research. Despite their promising performance, these models lack the ability to track object identities across multiple references or ground both actions and objects simultaneously.
To position GroundCap among the existing datasets, Table 1 presents a systematic comparison across key dimensions including scale, grounding capabilities, annotation types, and evaluation metrics.
Table 1: Comparison of GroundCap with related grounded captioning and visual grounding datasets. Avg Words refers to the average number of words per caption. Objects indicates whether the dataset includes object annotations. Actions indicates whether the dataset includes action annotations. Object Re-ID indicates whether the dataset can consistently track and ground all instances of the same object across multiple references within captions. Public refers to whether the full dataset is publicly available.
Dataset # Images Objects Actions Object Re-ID Avg Words Public
Flickr30k Entities [25] 31,783 -
Visual Genome [13] 108,249 5
DenseCap [14] 94,000 -
RefCOCO+ [10] 19,992 3.5
RefCOCOg [26] 25,799 8.4
KOSMOS-2 (GrIT) [3] 91,$0$
GROUNDHOG (M3G2) [4] 2,500,000 -
GLaMM (GranD-f) [24] 214,000 -
Groma Instruct [19] 30,000 -
Florence-2 (FLD-5B) [20] 126,$0$
GroundCap 52,016 128
@@ 3 The GroundCap Dataset
This section details the creation of GroundCap  [^1] The dataset is available for academic research under the same license as the underlying MovieNet dataset at https://huggingface.co/datasets/daniel3303/GroundCap. We first describe our frame selection process from MovieNet, followed by the object detection approach that handles both distinct objects and background elements. We then explain the caption generation pipeline that combines automated generation with human refinement.
@@ 3.2.2 Bounding Box Generation for “Stuff” Classes
Our approach to generate bounding boxes for “stuff” classes aims to provide a detailed representation of background elements while balancing with the lowest number of bounding boxes possible. The process begins by extracting pixel coordinates of the segmented area. K-means clustering is then applied to these coordinates, with K ranging from 1 to 6.
@@
For each clustering result, we calculate the 5th and 95th percentiles of the $x$ and $y$ coordinates within each cluster. These percentiles form the basis of our initial bounding boxes, excluding the 5% most extreme points on each end. This approach strikes a balance between coverage and compactness, ensuring that the initial boxes are representative of the cluster’s main body without being influenced by scattered outlier points.
@@
After generating these initial boxes, we perform an adjustment step to eliminate overlaps between boxes of the same class. Our adjustment algorithm iteratively identifies overlapping pairs of boxes and modifies their boundaries. When adjusting for overlaps, we prioritize minimal distortion by modifying the overlapping edges proportionally to the size of each box. We prefer adjusting the lowest overlapping axis with respect to the length of the boxes in the corresponding axis.
@@
For instance, we compare the horizontal overlap relative to the total width of the overlapping boxes against the vertical overlap relative to their total height. If the proportional horizontal overlap is smaller, we adjust the left and right edges; otherwise, we adjust the top and bottom edges, thus preserving the aspect ratio of the boxes as much as possible. The adjustment for each box is calculated based on its size relative to the total size of both overlapping boxes in the chosen dimension (i.e., larger boxes undergo larger adjustments). The process results in a set of non-overlapping boxes that adhere to the underlying clusters, balancing the need for coverage of the segmented area with minimal overflow into non-segmented regions and minimizing the number of bounding boxes. The final set of boxes is evaluated based on two primary criteria.
@@
Coverage: The proportion of the segmented area that is enclosed by the bounding boxes. We aim for a minimum coverage of 90% to ensure that the majority of the “stuff” region is represented. Overflow: The proportion of the area within the bounding boxes that does not belong to the segmented region. We aim for a maximum allowed overflow of 30% into non-segmented areas.
@@
The evaluation function balances these criteria, seeking to maximize coverage while minimizing overflow. This process is repeated 10 times for each number of clusters until a valid configuration is found. A configuration is considered valid when the coverage exceeds 90% and the overflow is below 30% or when the maximum number of clusters is reached. If multiple valid configurations are found for the same number of clusters, the best one is selected based on the configuration score calculated as the difference between coverage and overflow. This allows us to represent large or disconnected background elements with multiple focused bounding boxes.
@@ 3.2.3 Post-processing
We assign unique IDs to each detected object, following their spatial sorting order: e.g., person-0 will always appear before person-1 in the image, either in a higher band (closer to the top) or in the same band but further to the left. The consistent ordering ensures that object references in captions are predictable and spatially meaningful – when referring to “the person in the top left corner”, we can be confident it corresponds to person-0.
@@ 3.3 Grounded Caption Generation
The process begins by generating a general caption for the main image using Pixtral [17]. This initial caption provides a high-level description of the scene without explicit grounding information. The model receives the image along with a system prompt explaining the captioning task and a set of few-shot examples demonstrating proper caption format and style, following the few-shot learning paradigm established in recent work [16]. We then generate focused captions for each detected object. These object-specific captions are created by providing the model with cropped regions corresponding to the bounding box of each object. In the final stage, the LLM synthesizes a grounded caption by combining the main caption, object-specific captions, and detection information, such as object metadata, including class, ID, and coordinates of the bounding box. The model generates a grounded caption incorporating our grounding tag system, using three distinct tags: <gdo> for objects, <gda> for actions, and <gdl> for locations.
Quality control is implemented through an evaluation system based on precision ($P$), recall ($R$), and F1 score metrics. These values follow standard information retrieval definitions: $P=\text{TP}/(\text{TP}+\text{FP})$; $R=\text{TP}/(\text{TP}+\text{FN})$; and F1 as the harmonic mean of precision and recall. TP (true positives) represents correctly grounded objects, FP (false positives) are objects mentioned in the caption but not present in detections, and FN (false negatives) are detected objects omitted from the caption.
@@ 4 Grounded METEOR (gMETEOR)
5 Baseline Models
To establish a baselines, we fine-tuned Pixtral-12B [17] [^2] Fine-tuned model: https://huggingface.co/daniel3303/PixtralGroundCap. and Qwen2.5-VL 7B on GroundCap. Our baselines process the full image along with metadata about each detected object (including object ID, class, and normalized coordinates), enabling it to generate captions with appropriate grounding tags (<gdo>, <gda>, and <gdl>) that link text spans to specific objects in the image.